# Branch, or Layer? Zeroth-Order Optimization for Continual Learning of Vision-Language Models

### v1

@@ Abstract
Continual learning in vision-language models (VLMs) faces critical challenges in balancing parameter efficiency, memory consumption, and optimization stability. While First-Order (FO) optimization (e.g., SGD) dominate current approaches, their deterministic gradients often trap models in suboptimal local minima and incur substantial memory overhead. This paper pioneers a systematic exploration of Zeroth-Order (ZO) optimization for vision-language continual learning (VLCL). We first identify the incompatibility of naive full-ZO adoption in VLCL due to modality-specific instability. To resolve this, we selectively applying ZO to either vision or language modalities while retaining FO in the complementary branch. Furthermore, we develop a layer-wise optimization paradigm that interleaves ZO and FO across network layers, capitalizing on the heterogeneous learning dynamics of shallow versus deep representations. A key theoretical insight reveals that ZO perturbations in vision branches exhibit higher variance than language counterparts, prompting a gradient sign normalization mechanism with modality-specific perturbation constraints. Extensive experiments on four benchmarks demonstrate that our method achieves state-of-the-art performance, reducing memory consumption by 89.1% compared to baselines. Code will be available upon publication.
1 Introduction
@@
Continual Learning (CL) has witnessed significant advancements in convolutional architectures (e.g., ResNet erd ; res_2 ; res_3 ; pose and ViTvit_1 ; vit_2 ; vit_3 ). Recently, Vision-Language Models (VLMs)-based CL approaches have attracted growing research attention. Particularly, CLIP-based methods clip_cl_1 ; clip_cl_2 have demonstrated robust continual learning capabilities. For instance, ZSCL zscl introduces an unlabeled reference dataset for knowledge distillation to guide parameter space optimization. Mod-X mod-x maintains consistency in multimodal representation spaces through specially designed loss functions, effectively mitigating performance degradation caused by geometric variations in feature spaces. However, these methods require full-parameter fine-tuning of CLIP models, which incurs substantial computational overhead. The recently proposed MoE4Adapter MoE4Adapters addresses this limitation by adopting Parameter-Efficient Fine-Tuning (PEFT) strategies peft ; peft1 ; peft2 , achieving competitive CL performance with significantly reduced resource consumption.
Notably, existing VLM-based CL methods predominantly employ first-order (FO) optimization strategies (e.g., SGD, Adam) overview_opti . While these methods provide stable gradient descent directions, their inherent local convergence characteristics often lead to suboptimal solutions zo_intro_1 ; zo_intro_2 ; zo_intro_3 . MeZO mezo and MaZO maskzo both leverage ZO optimization to address the memory challenges of fine-tuning large language models in resource-constrained environments while maintaining or even surpassing the performance of traditional backpropagation-based fine-tuning. This breakthrough has inspired novel research directions in CL. For example, ZeroFlow zeroflow systematically compares various FO optimization strategies and discovers that ZO optimization exhibits unique advantages in alleviating catastrophic forgetting, providing new insights for optimizer design in continual learning systems.
However, the application of zero-order (ZO) optimization in Vision-Language Continual Learning (VLCL) remains understudied. Therefore, we conduct an in-depth exploration of ZO in VLCL using the state-of-the-art MoE4Adapter model MoE4Adapters as the baseline. First, we naively replace all FO optimizers with ZO counterparts to observe the direct impact of ZO in VLCL. However, experiments reveal that this approach causes excessive training divergence and failure to converge. Inspired by the cross-modal interaction consistency theory in VLMs cross_1 ; cross_2 , we introduce two investigations:
Investigation 1: Can ZO and FO collaborate for global model optimization across branches?
We implement ZO in either the visual or linguistic modality branch while retaining FO in the other. Experiments show that this strategy significantly outperforms the naive full-ZO approach.
Investigation 2: Can ZO and FO collaborate for hierarchical optimization within branch layers?
We thus deploy ZO only in specific layers (e.g., odd or consecutive layers) within a single branch while keeping FO in the remaining layers. Experiments further demonstrate that this approach yield better results, with certain configurations even surpassing the full-FO baseline. We attribute this to the stable directivity of FO optimization combined with the limited volatility of ZO, which helps the model escape local optima during optimization. Additionally, in layer-wise experiments, we find the visual branch exhibit significantly stronger fluctuations under ZO than the language branch. Therefore, we propose a stabilization mechanism into the visual branch to constrain ZO’s search range, further improving optimization convergence and overall continual learning performance.
@@
  • We reveal differential effects of ZO optimization on visual and language branches in VLCL and proposed a modality-specific framework combining ZO with FO optimization to balance stability and exploration.
  • Our layer-wise granular approach selectively deploys ZO in specific layers alongside FO in others, leveraging their synergy to escape local optima and enhance optimization efficiency.
  • Addressing ZO-induced volatility in the visual branch, we introduce a targeted stabilization mechanism to constrain its search range, significantly improving convergence stability and overall VLCL performance.
2 Optimization in Continual Learning: A Primer
Traditional optimization methods in continual learning primarily rely on FO gradient descent fo_1 ; fo_2 . Given a loss function $\mathcal{L}(\theta)$ and parameters $\theta$, FO updates parameters via:
$\theta=\theta-\eta\cdot\nabla_{\theta}\mathcal{L}(\theta),$ (1)
where $\nabla_{\theta}\mathcal{L}(\theta)$ is the FO gradient, and $\eta$ is the learning rate. While FO methods offer accurate gradient directions and have been widely adopted due to their efficiency, they may incur considerable memory overhead when higher-order derivatives (e.g., Hessians in second-order extensions) are involved lora ; memory_1 . Moreover, their deterministic update paths tend to limit exploration during training, potentially increasing the risk of convergence to suboptimal local minima and reducing adaptability in dynamic continual learning scenarios.
In contrast, ZO optimization estimates gradients through forward passes through purposeful perturbations. For a parameter subset $\theta^{k}_{m}$ (e.g., layer $k$ in modality $m$), ZO approximates gradients via directional perturbations:
@@ 2 Optimization in Continual Learning: A Primer
where $\Delta$ is a random directional vector, and $\varepsilon$ is a small perturbation scale. ZO methods avoid memory-intensive backward computations and introduce gradient stochasticity derivative , which may improve the ability to explore non-convex loss landscapes and escape poor local minima. However, the reliance on randomized perturbations can lead to variance in gradient estimates, whose effect may vary across different model architectures or modalities. These characteristics suggest that a hybrid optimization strategy, which balances FO’s directional precision with ZO’s exploratory potential, could offer a complementary pathway for improving training robustness in continual learning scenarios.
Figure 1: Illustration of our empirical study. The language encoder and vision encoder of CLIP are frozen, only the trainable units attached to each layer is performed to parameters update. To sum up, we systematically explores how ZO optimization operates in VLCL, including branches (Dual, Vision, or Language) and layers (w/ Hop-odd, w/ Hop-even, w/ Prefix (six) and w/ Suffix (six)).
3 Empirical Study of Zeroth-Order Optimization in VLCL
3.1 Implementation
@@
Datasets and task construction. We evaluate our method on three datasets: CIFAR-100 (100 fine-grained classes), Tiny-ImageNet (200-class subset of ImageNet), and ImageNet-R (45,000 artistic variants spanning 200 ImageNet classes). For task construction under the CIL paradigm, we adopt the IncX configuration, where a dataset is split into sequential tasks with $X$ classes per task (e.g., Inc20 denotes 5 tasks with 20 classes each on CIFAR-100). This configuration is consistently applied across all datasets. All tasks enforce disjoint class distributions and exclude task-specific identifiers during inference to ensure a rigorous evaluation protocol.
Method. We build on the MoE4Adapter MoE4Adapters , a state-of-the-art VLCL framework that incorporates MoE modules into pre-trained vision-language models (e.g., CLIP). To explore PEFT alternatives, we introduce a variant, MoE4Adapter†, which replaces the original MoE modules with single-layer Low-Rank Adaptation (LoRA lora ) modules. This modification streamlines the architecture by reducing computational overhead while preserving adaptability, allowing direct comparison between MoE and LoRA within a unified framework.
@@
Implementation details. All experiments employ the CLIP-ViT-B/16 backbone architecture, consistent with the baseline MoE4Adapter framework, to ensure direct comparability. Following MoE4Adapter’s paradigm, the pre-trained VLM backbone remains frozen, with only task-specific adapters (MoE or LoRA modules) being trainable. For the ZO-based method, we adopt a more conservative ZO strategy to enhance training stability and reduce the risk of noisy updates, which evaluates multiple candidate updates and selectively applies the one that yields the lowest loss. Hyperparameters including perturbation scale $\varepsilon=0.001$ for ZO gradients and FO/ZO mixing ratio $\lambda=1$, are validated on the first task and retained for subsequent tasks.
3.2 Rethinking Zeroth-Order Optimization in VLCL
The potential of zeroth-order optimization in VLCL. VLCL harnesses trainable units in VLMs to achieve parameter-efficient adaptation. Notably, trainable units often encounter optimization difficulties due to modality-specific feature distribution disparities. Conventional FO optimization method relies on precise gradient descent, hence leading to the attraction by local optima and suboptimal convergence. Besides, the necessity of back propagation requires storing high-gradients which significantly improve the memory overhead.
@@ 3.2 Rethinking Zeroth-Order Optimization in VLCL
To mitigate these challenges, we propose integrating ZO optimization into VLCL. By leveraging its perturbation-based search mechanism, ZO enhances exploration in the parameter space, enabling escape from local optima. Furthermore, ZO eliminates the need for storing high-order gradient information, significantly reducing memory overhead—a critical advantage for resource-constrained CL tasks. However, does naively replacing FO with ZO necessarily lead to better performance?
@@
Table 1: How ZO optimization affects VLCL through different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO optimization is applied. †indicates that the expert modules in MoE4Adapter are replaced with LoRA.
Method CIFAR-100 Inc20 CIFAR-100 Inc10 Tiny-Imagenet Inc20 Image-R Inc20
@@
Empirical analysis of zeroth-order optimization failure in VLCL. We attempt to apply ZO into the VLCL, a straightforward intuition is to replace FO optimizers with ZO methods across both the vision and language branches including all trainable units, and the results are shown in Table 1. However, it can be observed when ZO is adopted in both vision and language branch (dual w/ ZO), the performances are significantly degraded regardless of MoE or LoRA settings, with “Last.” decreasing by approximately 6%–11% and "Avg." by approximately 6%–13%. It can be attributed to the reason that the variance of ZO destabilizes VLCL training, leading to the optimization process is difficult to converge. To verify this point, we plot the loss function trajectories across these experimental settings shown in Figure 2. The Figure 2(a) and 2(b) reveal that dual-branch ZO approach suffers from severe loss oscillations and failed convergence compared to the original FO optimization. From this observation, we conclude that full ZO optimization is fundamentally ill-suited for VLCL, as its inherent gradient estimation fluctuations induce training instability. On the contrary, FO optimization provides stable gradient directions throughout the training process. This raises the question: Can synergistic integration of ZO and FO optimization achieve better performance?
@@
How can zeroth-order optimization be effective in VLCL? Branches, or Layers? We further investigate a heterogeneous optimization strategy: applying ZO to only one modality branch while retaining FO for the other. From a qualitative perspective, Figure 2(c) and 2(d) reveal when ZO is adopted into a single branch (vision or language), the overall loss function trajectory is promising to converge compared with dual branch ZO. We consider that the single-branch FO provides optimization stability, guiding the overall training process maintains consistency. Meanwhile, we can find that the performance obtains significant improvement shown in Table 1, most of the results are close to the baseline, and even some results outperform baseline. It can be explained that ZO’s perturbation-based gradient estimation introduces controlled stochasticity into the optimization process, which probabilistically assists the optimizer in evading suboptimal local minima. Additionally, we find that the performance of language w/ ZO generally outperform vision w/ ZO, we argue that the optimization stability of language branch is stronger than vision branch, which is verified by the loss trajectories of Figure 2(c) and 2(d). These findings suggest a ZO-FO synergy strategy which balances performance exploration and training convergence in VLCL, validating the feasibility of the integration optimization.
@@
3.3 Why Layers Matter: Triggering Effective Continual Adaptation
To explore the performance regarding layer-wise allocation of ZO and FO, four layer-wise ZO patterns were tested: Hop-odd (adopt ZO in odd layers), Hop-even (adopt ZO in even layers), Prefix (six) (adopt ZO in first six layers) and Suffix (six) (adopt ZO in last six layers) in dual or single modality branch, with FO used in other layers. The selection of modality branch is still refer to Table 1.
(a) Dual w/ Hop-odd.
(b) Vision w/ Hop-odd.
(c) Language w/ Hop-odd.
Figure 3: Analyzing convergence behavior of VLCL in Hop-odd across Dual, Vision, Language.
@@ 3.3 Why Layers Matter: Triggering Effective Continual Adaptation
Method CIFAR-100 Inc20 CIFAR-100 Inc10 Tiny-Imagenet Inc20 Image-R Inc20
@@
Layer-wise ZO unlocks VLCL performance potential. We then apply layer-wise ZO optimization to both dual and single modality branches, with the remaining layers optimized by FO. As shown in Table 2, we observe that layer-wise strategy can provide significant performance improvements compared to applying ZO across all trainable units in branches. In dual w/ ZO, the layer-wise ZO optimization averagely improves 9.4% accuracy on four patterns across all dataset. More interestingly, we observe that certain fine-grained layer-wise ZO patterns can outperform full FO approaches. To investigate the layer-wise effectiveness, we first analyze the performance of applying the collaborative ZO-FO optimization to a uniform layering strategy (with Hop-odd selected as a representative case). The loss trajectories under different modalities are recorded and presented in Figure 3, it demonstrates a more stable convergence process during training compared to that shown in Figure 2. We argue that more fine-grained layer-wise strategy further amplifies the respective advantages of ZO and FO in VLCL, FO provides stable gradient directions and ZO’s stochastic perturbations help the optimization process escape local minima.
@@
3.4 A New Enhancement from Vision Branch Discrepancy

### v2

@@ Abstract
Vision-Language Continual Learning (VLCL) has attracted significant research attention for its robust capabilities, and the adoption of Parameter-Efficient Fine-Tuning (PEFT) strategies is enabling these models to achieve competitive performance with substantially reduced resource consumption. However, dominated First-Order (FO) optimization is prone to trap models in suboptimal local minima, especially in limited exploration subspace within PEFT. To overcome this challenge, this paper pioneers a systematic exploration of adopting Zeroth-Order (ZO) optimization for PEFT-based VLCL. We first identify the incompatibility of naive full-ZO adoption in VLCL due to optimization process instability. We then investigate the application of ZO optimization from a modality branch-wise to a fine-grained layer-wise across various training units to identify an optimal strategy. Besides, a key theoretical insight reveals that vision modality exhibit higher variance than language counterparts in VLCL during the ZO optimization process, and we propose a modality-aware ZO strategy, which adopts gradient sign normalization in ZO and constrains vision modality perturbation to further improve performance. Benefiting from the adoption of ZO optimization, PEFT-based VLCL fulfills better ability to escape local minima during the optimization process, extensive experiments on four benchmarks demonstrate that our method achieves state-of-the-art results.
Introduction
@@
Continual Learning (CL) has witnessed significant advancements in convolutional architectures (e.g., ResNet (erd; res_2; pose) and ViT(vit_1; vit_2; vit_3)). Recently, Vision-Language Models-based Continual Learning (VLCL) approaches have attracted growing research attention. Particularly, CLIP-based methods (clip_cl2; lwf-vr; zhao2; mod-x) have demonstrated robust continual learning capabilities. However, these methods require full-parameter fine-tuning of CLIP models, which incurs substantial computational overhead. To overcome this critical bottleneck, Parameter-Efficient Fine-Tuning (PEFT) strategies (peft; peft1; peft2; dmnsp) have recently emerged as a compelling alternative. These techniques make it possible to achieve competitive CL performance with significantly reduced resource consumption. For instance, the VLCL method MoE4Adapter (MoE4Adapters) leverages a PEFT approach to effectively address this limitation.
Existing VLCL methods predominantly employ First-Order (FO) optimization strategy (overview_opti), which update parameters using precise gradients derived from backpropagation. While valued for their stable directional guidance, this approach becomes a drawback in the context of PEFT. Its deterministic update paths limit exploration during training , and the low-dimensional subspace to which PEFT confines optimization makes these methods susceptible to converging to sharp local optima that overfit the current task (SharpMinima; sam), leading to a performance drop when faced with new tasks and potentially exacerbating catastrophic forgetting. To explore solutions with stronger generalization capabilities, Zeroth-Order (ZO) optimization offers a promising alternative (zeroflow). Unlike traditional FO optimization, this method forgoes precise gradients from backpropagation, instead estimating performance with random perturbations, making it less prone to getting trapped in local minima when exploring a constrained space (mezo; maskzo), hence potentially applied in PEFT-based VLCL.
There is already a vast amount of ZO finetuning researches (mezo; maskzo), while they lacks of considering its common strategy of fully replacing FO optimization is applicable to VLCL, given the documented optimization disparities between modality branches (vlgap; clap4clip; vldis1; vldis2; vldis3) and the inherent sensitivity of ZO’s perturbation-based approach. In this paper, we initiate the study of applying ZO optimization into VLCL, and aim to answer the following key question: How can ZO be integrated in VLCL settings, and can it boost overall performance?
To answer the question, this paper adopts a fine-grained perspective, investigating the performance of ZO when applied to different modality branches and trainable layers within PEFT-based VLCL. Specifically, we respectively explore ZO in the vision or language modality branch while retaining FO in the other, to identify their benefits and limitations. Based on the insights gained, we then extend our investigation to a more granular layer-wise and adopt ZO into different training units, including continuous prefix/suffix layers and interleaved odd/even layers across a modality branch, to obtain a optimal result. Meanwhile, we identify a convergence discrepancy between modalities of VLCL under ZO optimization, thus proposing a Modality-aware ZO (MoZO) strategy, adopting gradient sign normalization in ZO and constraining vision modality perturbation to further boost VLCL performance.
Figure 1: Illustration of our study. The language and vision encoders of CLIP are frozen, only the trainable units attached to each layer is performed to parameters update. To sum up, we systematically explores how ZO optimization operates in VLCL, including branches (Dual, Vision, or Language) and layers (w/ Hop-odd, w/ Hop-even, w/ Prefix (six) and w/ Suffix (six)).
@@
  • We present the optimization challenge in PEFT-based VLCL, where conventional FO method is prone to converging to suboptimal local optima, and explore to leverage the ZO optimization to address the problem.
  • We adopt the ZO optimization at both branch-wise and fine-grained layer-wise within PEFT-based VLCL, achieving optimal performance through detailed analysis and refined application strategies.
  • We identify the issue of optimization discrepancy between modalities of PEFT-based VLCL under ZO optimization, and propose a MoZO strategy further improving overall performance.
A Preliminary for ZO Optimization in PEFT-based VLCL
Traditional optimization methods in continual learning primarily rely on FO gradient descent (fo_1; fo_2), updating model parameters $\theta$ based on precise gradients computed via backpropagation:
$\theta_{t+1}=\theta_{t}-\eta_{t}\cdot\nabla_{\theta}\mathcal{L}(\theta_{t}),$ (1)
where $\nabla_{\theta}\mathcal{L}(\theta_{t})$ is the first-order gradient. While FO methods offer accurate gradient directions and have been widely, their deterministic update paths tend to limit exploration during training, potentially increasing the risk of convergence to suboptimal local minima and reducing adaptability in dynamic continual learning scenarios.
In contrast, ZO optimization estimates gradients through forward passes through purposeful perturbations. For a parameter subset $\theta^{k}_{m}$, ZO approximates gradients via directional perturbations:
@@ A Preliminary for ZO Optimization in PEFT-based VLCL
where $\Delta$ is a random directional vector, and $\varepsilon$ is a small perturbation scale. ZO methods introduce gradient stochasticity (derivative), which may improve the ability to explore non-convex loss landscapes and escape poor local minima. However, the reliance on randomized perturbations can lead to variance in gradient estimates, whose effect may vary across different model architectures or modalities. To validate this point, we first establish a baseline and investigate the most straightforward strategy: a complete replacement of the FO optimizer with ZO in PEFT-based VLCL, which means adopt ZO in both vision-language branches and overall trainable units. However, our subsequent experiments demonstrate this strategy leads to severe training instability, evidenced by loss oscillations during the convergence process, and results in a significant performance degradation.
Recent research suggests that partially incorporating exploratory while high-variance ZO estimators into the model architecture improves its global optimization performance (hybridzo1; hybridzo2). We translate this advance to the context of PEFT-based VLCL. As shown in Figure 1, we adopt a empirical exploration to progressively investigate how to best apply the ZO optimization in VLCL. To explore suitable application paradigms, we designed a comprehensive set of experiments applying ZO at varying levels of granularity, from entire modality branches to fine-grained partially trainable units. However, given the vast diversity of configurations arising from combinations of modality branches and PEFT trainable units, constructing a single, unified theoretical framework is intractable. Consequently, we conduct a comprehensive empirical analysis to investigate both the differential impacts of applying ZO across modality branches and the distinct characteristics of employing it in various trainable layer configurations (such as continuous and interleaved layers), which allows us to achieve optimal results.
Study of ZO Optimization in VLCL
Implementation
@@
Datasets and task construction. We evaluate our method on three datasets: CIFAR-100 (CIFAR), Tiny-ImageNet (TinyImg), and ImageNet-R (ImgR). For task construction under the CIL paradigm, we adopt the IncX configuration (e.g., Inc20 denotes 5 tasks with 20 classes each on CIFAR). All tasks enforce disjoint class distributions and exclude task-specific identifiers during inference to ensure a rigorous evaluation protocol.
Baseline. We choose MoE4Adapter (MoE4Adapters) as the SOTA baseline, which incorporates Mixture of Experts (MoE) into CLIP for VLCL. To explore PEFT alternatives, we also replaces the MoE modules with Low-Rank Adaptation (LoRA (lora)) modules. For LoRA results, we present a part of them and put the remaining in supplementary material.
@@
Implementation details. All experiments employ the CLIP-ViT-B/16 backbone architecture. Following baseline’s paradigm, the CLIP backbone remains frozen, with only task-specific adapters (MoE or LoRA modules) being trainable. For the ZO-based method, we adopt a more conservative ZO strategy which evaluates multiple candidate updates and selectively applies the one that yields the lowest loss(zeroflow). Hyperparameters including perturbation scale $\varepsilon=0.001$ for ZO gradients and FO/ZO mixing ratio $\lambda=1$, are validated on the first task and retained for subsequent tasks.
Rethinking ZO Optimization in VLCL
@@
The potential of ZO optimization in VLCL. PEFT-based VLCL harnesses trainable units in vision-language model to achieve parameter-efficient adaptation. Conventional FO optimization method relies on precise gradient descent, hence leading to the attraction by local optima and suboptimal convergence. To mitigate these challenges, we first explore integrate full ZO optimization into VLCL. By leveraging its perturbation-based search mechanism, ZO enhances exploration in the parameter space, enabling escape from local optima. However, does naively replacing FO with ZO necessarily lead to better performance?
@@
Table 1: How ZO optimization affects VLCL through different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO optimization is applied. The $\dagger$} indicates that the MoE modules in baseline are replaced with LoRA.
Method CIFAR Inc20 CIFAR Inc10 TinyImg Inc20 ImgR Inc20
@@
Analysis of naive ZO optimization failure in VLCL. We attempt to apply ZO into the VLCL, a straightforward intuition is to replace FO optimizers with ZO methods across both the vision and language branches including all trainable units, and the results are shown in Table 1. However, it can be observed when ZO is adopted in both vision and language branch (dual w/ ZO), the performances are significantly degraded regardless of MoE or LoRA settings, with Last. and Avg. averagely decreasing by 8.5% and 9.5% respectively. It can be attributed to the reason that the variance of ZO destabilizes VLCL training, leading to the optimization process is difficult to converge. To verify this point, we plot the loss function trajectories across these experimental settings shown in Figure 2. The Figure 2(a) and 2(b) reveal that dual-branch ZO approach suffers from severe loss oscillations and failed convergence compared to the original FO optimization. From this observation, we conclude that full ZO optimization is fundamentally ill-suited for VLCL, as its inherent gradient estimation fluctuations induce training instability. On the contrary, FO optimization provides stable gradient directions throughout the training process. This raises the question: Can synergistic integration of ZO and FO optimization achieve better performance?
@@
How can ZO optimization be effective in VLCL? Branches, or Layers? We further investigate a heterogeneous optimization strategy: applying ZO to only one modality branch while retaining FO for the other. From a qualitative perspective, Figure 2(c) and 2(d) reveal when ZO is adopted into a single branch (vision or language), the overall loss function trajectory is promising to converge compared with dual branch ZO. We consider that the single-branch FO provides optimization stability, guiding the overall training process maintains consistency. Meanwhile, we can find that the performance obtains significant improvement shown in Table 1, most of the results are close to the baseline, and even some results outperform baseline. It can be explained that ZO’s perturbation-based gradient estimation introduces controlled stochasticity into the optimization process, which probabilistically assists the optimizer in evading suboptimal local minima. Additionally, we find that the performance of language w/ ZO generally outperform vision w/ ZO, we argue that the optimization stability of language branch is stronger than vision branch, which is verified by the loss trajectories of Figure 2(c) and 2(d). These findings suggest a ZO-FO synergy strategy which balances performance exploration and training convergence in VLCL, validating the feasibility of the integration optimization.
@@
Why Layers Matter: Triggering Effective Continual Adaptation
To explore the performance regarding layer-wise allocation of ZO and FO, four layer-wise ZO patterns were tested: Hop-odd (adopt ZO in odd layers), Hop-even (adopt ZO in even layers), Prefix (six) (adopt ZO in first six layers) and Suffix (six) (adopt ZO in last six layers) in dual or single modality branch, with FO used in other layers. The baseline and selection of modality branch is still refer to Table 1.
(a) Du. w/Hop-odd.
(b) Vis. w/Hop-odd.
(c) Lan. w/Hop-odd.
Figure 3: Analyzing convergence behavior of VLCL in Hop-odd across Dual (Du.), Vision (Vis.), Language (Lan.).
@@
Method CIFAR Inc20 CIFAR Inc10 TinyImg Inc20 ImgR Inc20
@@
Layer-wise ZO unlocks VLCL performance potential. We then apply layer-wise ZO optimization to both dual and single modality branches, with the remaining layers optimized by FO. As shown in Table 2, we observe that layer-wise strategy can provide significant performance improvements compared to applying ZO across all trainable units in branches. In dual w/ ZO, the layer-wise ZO optimization averagely improves 9.4% accuracy on four patterns across all dataset. More interestingly, we observe that certain fine-grained layer-wise ZO patterns can outperform full FO approaches. To investigate the layer-wise effectiveness, we first analyze the performance of applying the collaborative ZO-FO optimization to a uniform layering strategy (with Hop-odd selected as a representative case). The loss trajectories under different modalities are recorded and presented in Figure 3, it demonstrates a more stable convergence process during training compared to that shown in Figure 2. We argue that more fine-grained layer-wise strategy further amplifies the respective advantages of ZO and FO in VLCL, FO provides stable gradient directions and ZO’s stochastic perturbations help the optimization escape from local minima.
@@
A New Enhancement from Vision Discrepancy